10. Daubert, Inc, planned to issue and sell at par 10-year, $1,000 face value bonds totaling $400 million next month. The bonds have been printed with a 6% coupon rate. Since that printing, however, Moody’s downgraded Daubert’s bond rating from Aaa to Aa. This means the bonds will have to offered to yield buyers 7%. How much less than it expected will Daubert collect when the bonds are issued?
Answer #1

Using financial calculator
Input: FV= 1000

PMT=6%*1000=60

N=10

I/Y = 7

Solve for PV as -929.76

Hence amount collected = 400000,000*929.76/1000 =

371904000

Difference in amount collected =

28096000